The probability that a new advertising campaign will increase sales is assessed as being 0.75. The probability that the cost of developing the new ad campaign can be kept within the original budget allocation is 0.60. Assuming that the two events are independent, the probability that the cost is not kept within budget or the campaign will not increase sales is ________. (Round your answer to three decimal places. Express your answers as a probability, not a percentage.)
Answer #1

A = Cost kept in budget,

P(A) = 0.75 , So P(A') =1 - P(A) = 1 - 0.75 = 0.25  

B = campaign will increase sales

P(B) = 0.60 So, P(B') = 1 - P(B) = 1 - 0.60 = 0.40

So,

P(A' B') = P(A') + P(B') - P(A' B')

= P(A') + P(B') - P(A' ) * P(B') [ Since A and B independent, P(A' B') = P(A' ) * P(B') ]

= 0.25 + 0.40 - 0.25 * 0.40

= 0.55